OpenStudy (anonymous):

solve each equation: 1/x = 5/x-4

OpenStudy (anonymous):

cross multiply 5x=x-4 4x=-4 x=-1

OpenStudy (anonymous):

what about 2/x+3 - 1/x= -6/x(x+3)

OpenStudy (anonymous):

holy crap what grade math is this?

OpenStudy (anonymous):

it's algebra 2

OpenStudy (anonymous):

well ican tell u x(x+3) is \[x^{2}\]+ 3x

OpenStudy (anonymous):

srry im no help im in geometry

OpenStudy (anonymous):

but you got the first problem right though

OpenStudy (anonymous):

lol 2nd is a lot harder

OpenStudy (anonymous):

Well, i have finals on thursday. It's very sad when a geometry student knows more algebra than me

OpenStudy (anonymous):

id say move everything to 1 side and set it all to equal to 0 and see how that works. if that doesnt then just cross multiply the one side into both fractions

OpenStudy (anonymous):

lol ur not doing bad. and im doing my review for geometry right now too. my finals r june 15th though

OpenStudy (anonymous):

school ends on june 8th for me

OpenStudy (anonymous):

same here